Find the pH of a formic acid solution that was is prepared by adding 0.280 g of formic acid to a volumetric flask and bringing the final volume to 1.44 L by adding water
Answer #1

molarity of HCOOH = W/G.M.Wt * volume of solution in L

                                  = 0.280/(46*1.44)

                                   = 0.00423M

               HCOOH(aq) + H2O(l) -----------------> HCOO^-(aq) + H3O^+ (aq)

I              0.00423                                                  0                     0

C                  -x                                                      +x                    +x

E           0.00423-x                                                 +x                     +x

                    Ka   =   [HCOO^-][H3O^+]/[HCOOH]

                   1.8*10^-4 = x*x/(0.00423-x)

                   1.8*10^-4*(0.00423-x) = x^2

                       x = 0.0007872

               [H3O^+]   =x   = 0.0007872M

                PH   = -log[H3O^+]

                          = -log0.0007872

                          = 3.104 >>>>answer